# Nonparametric Jackknife Instrumental Variable Estimation and Confounding Robust Surrogate Indices

### v1

@@ 1 Introduction
    In Section 4, we a develop methods for inference on the new experiment’s long-term causal effect. We devise a debiased estimator that involves a new debiasing nuisance related to the ratio of densities of short-term outcomes in the old and the new experiments. We develop an estimator for this debiasing nuisance with guarantees in the many-weak-IV setting. And, we show that when we combine npJIVE with this debiasing nuisance in the right way, we can obtain asymptotically normal estimates of the long-term causal effect in the new experiment, even when npJIVE and the debiasing nuisance are learned completely nonparametrically. This ensures fast estimation rates and implies simple Wald confidence intervals give asymptotically correct coverage.
@@ 2 The npJIVE Estimator for NPIV with Many Weak IVs
We now motivate and define npJIVE. For function $f(A,S,Y)$ let us write $[\Pi f](A)=\mathbb{E}[f(A,S,Y)\mid A]$ and $\|f\|^{2}=\frac{1}{K}\sum_{a=1}^{K}\|f\|^{2}_{(a)}$ with $\|f\|^{2}_{(a)}=\mathbb{E}[f^{2}(a,S,Y)\mid A=a]$, and let $L_{2}(S)=\{h(S):\|h\|<\infty\}$ be the square integrable functions of $S$. Let
$\mathcal{H}_{0}=\{h\in L_{2}(S):R_{1}(h)=0\},~{}~{}\text{where}~{}~{}R_{1}(h)=\frac{1}{2}\|\Pi(h-Y)\|^{2},$
be the solutions to Eq. 1. Given a hypothesis class $\mathcal{H}\subseteq L_{2}(S)$, suppose $\mathcal{H}$ contains a solution $h_{0}\in\mathcal{H}\cap\mathcal{H}_{0}\neq\varnothing$. Then $h_{0}\in\operatornamewithlimits{argmin}_{h\in\mathcal{H}}R_{1}(h)\subseteq\mathcal{H}_{0}$, meaning the solution must minimize the risk criterion $R_{1}(h)$ over $\mathcal{H}$ and any such minimizer must solve Eq. 1.
@@
$\widehat{R}_{1}^{\mathrm{plug-in}}(h)=\frac{1}{2K}\sum_{a=1}^{K}([\widehat{\Pi}(Y-h)](a))^{2},~{}~{}\text{where}~{}~{}[\widehat{\Pi}f](a)=\frac{1}{n}\sum_{i:A_{i}=a}f(a,S_{i},Y_{i}),$ (2)
meaning $[\widehat{\Pi}f](a)$ is the sample average over the $A_{i}=a$ data. In fact, optimizing $\widehat{R}_{1}^{\mathrm{plug-in}}(h)$ over linear functions $h(S)=S^{\top}\beta$ is almost 2SLS (aside from some $A$-dependent sample weights). And, optimizing it over a nonparametric hypothesis class $\mathcal{H}$ (possibly with regularization) is exactly the adversarial NPIV estimators of (Bennett et al., 2019, 2023a, 2023b; Bennett and Kallus, 2023; Dikkala et al., 2020) with an adversary function class that includes all functions of the discrete $A$.
@@
$\widehat{R}_{1}(h)$ $=\frac{1}{2K}\sum_{a=1}^{K}[\widehat{\Pi}_{0}(Y-h)](a)[\widehat{\Pi}_{1}(Y-h)](a),$
$\text{where}~{}~{}[\widehat{\Pi}_{v}f](a)$ $=\frac{2}{n}\sum_{i:A_{i}=a,V_{i}=v}f(a,S_{i},Y_{i}),$
assuming $n$ is even. Meaning $[\widehat{\Pi}_{v}f](a)$ takes the sample average over the observations for which $A_{i}=a$ belonging to fold $v$.
@@ 2.1 Analysis of npJIVE
The relevant function classes for analyzing npJIVE are: $\mathcal{H}^{0}=\mathcal{H}-h_{0}$, $\mathcal{G}^{0}=\{g(S)-(\Pi g)(A):g\in\mathcal{H}^{0}\}$, $\mathcal{G}^{1}=\{g(S)(\Pi g)(A):g\in\mathcal{H}^{0}\}$, and $\widehat{\mathcal{G}}=\{g(S)(\widehat{\Pi}_{0}g)(A):g\in\mathcal{G}^{0}\}$.
Let $\mathrm{star}(\mathcal{F})=\{\gamma g:\gamma\in[0,1],g\in\mathcal{F}\}$. Let $\widetilde{\delta}_{N}$, $\bar{\delta}_{N}$ and $\breve{\delta}_{N}$ be solutions to $\mathcal{R}_{N}(\mathrm{star}(\mathcal{F}),\delta)\leq\delta^{2}$ for $\mathcal{F}=\mathcal{G}^{0},\mathcal{G}^{1},\Pi\mathcal{G}$, respectively, let $\widetilde{\delta}_{n}$ be a solution to $\mathcal{R}_{n,a}(\mathrm{star}(\mathcal{G}^{0}),\delta)\leq\delta^{2}$ for every $a$, and let $\widehat{\delta}_{N}$ be a solution to $\mathcal{R}_{N}^{|0}(\mathrm{star}(\widehat{\mathcal{G}}),\delta)\leq\delta^{2}$. Let $\widehat{r}_{N}^{2}$ and $\bar{r}_{N}^{2}$ be bounds on $\mathcal{R}^{|0}(\widehat{\mathcal{G}},\infty)$ and $\mathcal{R}(\mathcal{G}^{1},\infty)$.
@@
$\|\Pi(\widehat{h}-h_{0})\|^{2}=O(\lambda+\delta_{N}\delta_{n})~{}~{}\text{with probability $1-O(\zeta)$, and}$
@@
3 Confounding Robust Surrogate Indices from Many Weak Experiments
@@ 3.1 Background: Statistical-Surrogate Indices
Possibly the most common approach assumes that short-term observations form a statistical surrogate (Prentice, 1989). This assumes that the long-term outcome is conditionally independent of treatment given the short-term outcomes. This involves two key restrictions: that there are no unobserved confounding between short- and long-term outcomes (no $U_{1}$ in Fig. 1(a)) and that all of the treatment’s effect on long-term outcomes is mediated by the short term (no $U_{2}$ in Fig. 1(a)). As the latter restriction becomes more defensible as we include more short-term outcomes so as to mediate more of the treatment’s long-term effect, Athey et al. (2019) combine many short-term observations into a surrogate index, assuming they form a statistical surrogate and using historical data to regress long-term on short-term (or other ways of adjustment such as weighting). The proposal is simple, effective, and, as such, widely adopted.
However, even if short-term outcomes fully mediate the long-term effect, they may fail to satisfy statistical surrogacy. Consider the causal diagram in Fig. 1(a), where $S$ perfectly mediates $A$’s effect on $Y$ (i.e., exclusion restriction), but $S$ and $Y$ share an unobserved confounder $U$, while treatment $A$ is fully unconfounded. In this case, $S$ is a collider so that conditioning on it induces a path from $A$ to $Y$ via $U$, violating surrogacy and imperiling analysis using such methods as Athey et al. (2019).
@@ 3.2 Experiments as Instruments and Surrogates as Proxies
Athey et al. (2019) consider a setting where historical data prior to the present experiment contains only $S$ and $Y$ and some baseline covariates. In this context, all we can do is either worry about potential unobserved confounders, or hope the included covariates satisfy ignorability between $S$ and $Y$. However, at organizations that routinely run many digital experiments, we can take historical data from past experiments where we also observe the randomized treatments $A$. In the setting of Fig. 1(a), these treatments constitute an IV, which can help us identify the causal effect of $S$ on $Y$ and therefore infer the effect of a novel treatment on $Y$ by considering only its effect on $S$.
@@
In the following, we first present the corresponding causal model (summarized in Fig. 1(b)). Then, we present an identification result showing how the long-term outcome of a novel treatment can be identified without observations on long-term outcomes under the novel treatment by leveraging historical experiments where we observe both short- and long-term outcomes. This identification result is written as a function of NPIV solutions, which may be estimated slowly in this nonparametric setting, imperiling rates on causal effect estimation if we were to simply plug-in an estimated solution. Therefore, we finally present a debiased formulation of the identification result that, under a strong identification condition, enjoys a mixed-bias property, which means we can multiply estimation rates and make them insignificant to first order. Together with estimation results for a nuisance in this new debiased formulation, we obtain asymptotic normal estimates for long-term causal effects of the novel treatment, enabling inference.
@@
(a) A setting with unconfounded treatment ($A$) but confounded surrogate ($S$) and outcome ($Y$).
[…]
(b) The general setting we tackle with two kinds of possible short-term observations $S=(S_{1},S_{2})$. The identity of the components $S_{1}$ and $S_{2}$ within $S$ need not be known.
@@ 3.3 Causal Model
We state our causal model in general terms using potential outcomes. The model is summarized in Fig. 1(b), which represents a causal diagram that is consistent with these assumptions. We assume the existence of variables $\widetilde{A},U_{1},S_{1}(a),U_{2}(a),S_{2}(u_{2}),Y(s_{1},s_{2},u_{2})$, and we abbreviate $S_{2}(a)=S_{2}(U_{2}(a))$, $Y(a)=Y(S_{1}(a),S_{2}(a),U_{2}(a))$.
Here we let $\widetilde{A}\in\{1,\dots,K+1\}$, where $\widetilde{A}\leq K$ denotes the historical experimental data and $\widetilde{A}=K+1$ denotes the novel treatment (we will return to data generation after the causal model). We are interested in inference on the average outcome under the novel treatment:
$\theta_{0}=\mathbb{E}[Y(K+1)].$
@@
$A$ is independent of all potential outcomes and of $U_{1}$.
Assumption 3 (Existence of bridge function).
There exists $h^{\star}$ such that, for every $a\in[K]$, $\mathbb{E}[Y(a)-h^{\star}(S(a))]=0$.
Assumption 4 (Extrapolability).
For any $h:\mathcal{S}\to\mathbb{R}$
$(\mathbb{E}[Y(a)-h(S(a))]~{}\forall a\in[K])\implies\mathbb{E}[Y(K+1)-h(S(K+1))]=0.$ (6)
@@
It is straightforward to check that our causal assumptions hold under the following nonparametric structural causal model (SCM)
$Y=$ $f_{Y}(S_{1},S_{2},U_{2},\epsilon_{Y})+g_{Y}(U_{1},S_{2},U_{2},\epsilon_{Y})$
$S_{2}=$ $f_{S_{2}}(U_{2},\epsilon_{S_{2}})$
$S_{1}=$ $f_{S_{1}}(A,U_{1},\epsilon_{S_{1}})$
$U_{2}=$ $f_{U_{2}}(A,\epsilon_{U_{2}})$
$U_{1}=$ $f_{U_{1}}(\epsilon_{U_{1}})$
where $\epsilon_{Y}$, $\epsilon_{S_{2}}$, $\epsilon_{S_{1}}$, $\epsilon_{U_{2}}$, $\epsilon_{U_{1}}$ are independent. The key for assumption 3 to hold is that $U_{1}$ and $S_{1}$ enter the outcome equation in an additively separable manner. Aside from this additive restriction, this particular SCM is depicted in Fig. 1(b).
@@ 3.4 Data
Let $\widetilde{S}=(S_{1}(\widetilde{A}),S_{2}(\widetilde{A}))$ and $\widetilde{Y}=Y(\widetilde{A})$ be the factually observed outcomes.
@@
Let $S^{(K+1)}\sim\widetilde{S}\mid\widetilde{A}=K+1$. We observe $n^{\prime}$ independent copies $S^{(K+1)}_{1},\ldots,S^{(K+1)}_{n^{\prime}}$ of $S^{(K+1)}$.
We now turn to the question of identification. Note that had we observed $Y^{(K+1)}\sim\widetilde{Y}\mid\widetilde{A}=K+1$ from the novel treatment, we would have trivially identified $\theta_{0}$ as $\mathbb{E}[Y^{(K+1)}]$ under Assumption 2. So, in the absence of outcome observations for the novel treatment, our goal is to combine the $N$ and $n^{\prime}$ datasets to identify $\theta_{0}$.
@@ 3.5 Surrogacy and set identification
It is immediate that under assumptions 3 and 4, $h^{\star}$ is a valid surrogate function in the sense that $E[Y(K+1)]=E[h^{\star}(S(K+1))]$.
That is, $h^{\star}$ connects causal effects on $Y$ to causal effects on $S$: if we could access $h^{\star}$, we would be able to know the average potential outcome of $Y$ under any treatment from the distribution of the average potential outcome of $S$ under that same intervention. We thus define the causal parameter of interest as $\theta^{\star}=\theta(h^{\star})$ where $\theta(h)=\mathbb{E}[h(S(a^{\prime}))]$ for the novel intervention $a^{\prime}$.
@@
For any function $h:\mathcal{S}\to\mathbb{R}$, let $[\Pi h](a)=\mathbb{E}[h(S)\mid A=a]$. The following lemma characterizes $h^{\star}$ in terms of the data distribution in the historical cells.
@@
Under the causal assumptions 1-3, it holds that $h^{\star}\in\mathcal{H}_{0}$ where $\mathcal{H}_{0}$ is the set of functions $h:\mathcal{S}\to\mathbb{R}$ that satisfy the conditional moment restriction (CMR) $[\Pi h](a)=\mathbb{E}[Y\mid A=a]$ for every $a\in[K]$.
3.6 Point identification and strong identification
While $h^{\star}$ is not point identified a priori, it is not our parameter of interest: $\theta(h^{\star})$ is. We now present some assumptions under which $\theta(h^{\star})$ is point identified.

### v2

@@ 1 Introduction
    In Section 4.1, we a develop methods for inference on the new experiment’s long-term causal effect. We devise a debiased estimator that involves a new debiasing nuisance related to the ratio of densities of short-term outcomes in the old and the new experiments. We develop an estimator for this debiasing nuisance with guarantees in the many-weak-IV setting. And, we show that when we combine npJIVE with this debiasing nuisance in the right way, we can obtain asymptotically normal estimates of the long-term causal effect in the new experiment, even when npJIVE and the debiasing nuisance are learned completely nonparametrically. This ensures fast estimation rates and implies simple Wald confidence intervals give asymptotically correct coverage.
@@ 2 The npJIVE Estimator for NPIV with Many Weak IVs
We now motivate and define npJIVE. For function $f(A,S,Y)$ let us write $[\mathcal{T}_{K}f](A)=\mathbb{E}[f(A,S,Y)\mid A]$and $\|f\|^{2}=\frac{1}{K}\sum_{a=1}^{K}\|f\|^{2}_{(a)}$ with $\|f\|^{2}_{(a)}=\mathbb{E}[f^{2}(a,S,Y)\mid A=a]$, and let $L_{2}(S)=\{h(S):\|h\|<\infty\}$ be the square integrable functions of $S$.
Let
$\mathcal{H}_{0}=\{h\in L_{2}(S):R_{1}(h)=0\},~{}~{}\text{where}~{}~{}R_{1}(h)=\frac{1}{2}\|\mathcal{T}_{K}(h-Y)\|^{2},$
be the solutions to Eq. 1.
Given a hypothesis class $\mathcal{H}\subseteq L_{2}(S)$, suppose $\mathcal{H}$ contains a solution $h_{0}\in\mathcal{H}\cap\mathcal{H}_{0}\neq\varnothing$. Then $h_{0}\in\operatornamewithlimits{argmin}_{h\in\mathcal{H}}R_{1}(h)\subseteq\mathcal{H}_{0}$, meaning the solution must minimize the risk criterion $R_{1}(h)$ over $\mathcal{H}$ and any such minimizer must solve Eq. 1.
@@
$\widehat{R}_{1}^{\mathrm{plug-in}}(h)=\frac{1}{2K}\sum_{a=1}^{K}([\widehat{\mathcal{T}}_{K}(Y-h)](a))^{2},~{}~{}\text{where}~{}~{}[\widehat{\mathcal{T}}_{K}f](a)=\frac{1}{n}\sum_{i:A_{i}=a}f(a,S_{i},Y_{i}),$ (2)
meaning $[\widehat{\mathcal{T}}_{K}f](a)$ is the sample average over the $A_{i}=a$ data. In fact, optimizing $\widehat{R}_{1}^{\mathrm{plug-in}}(h)$ over linear functions $h(S)=S^{\top}\beta$ is almost 2SLS (aside from some $A$-dependent sample weights). And, optimizing it over a nonparametric hypothesis class $\mathcal{H}$ (possibly with regularization) is exactly the adversarial NPIV estimators of (Bennett et al., 2019, 2023a, 2023b; Bennett and Kallus, 2023; Dikkala et al., 2020) with an adversary function class that includes all functions of the discrete $A$.
@@
$\widehat{R}_{1}(h)$ $=\frac{1}{2K}\sum_{a=1}^{K}[\widehat{\mathcal{T}}_{K,0}(Y-h)](a)[\widehat{\mathcal{T}}_{K_{1}}(Y-h)](a),$
$\text{where}~{}~{}[\widehat{\mathcal{T}}_{K,v}f](a)$ $=\frac{2}{n}\sum_{i:A_{i}=a,V_{i}=v}f(a,S_{i},Y_{i}),$
assuming $n$ is even. Meaning $[\widehat{\mathcal{T}}_{K,v}f](a)$ takes the sample average over the observations for which $A_{i}=a$ belonging to fold $v$.
@@ 2.1 Analysis of npJIVE
The relevant function classes for analyzing npJIVE are: $\mathcal{H}^{0}=\mathcal{H}-h_{0}$, $\mathcal{G}^{0}=\{g(S)-(\mathcal{T}_{K}g)(A):g\in\mathcal{H}^{0}\}$, $\mathcal{G}^{1}=\{g(S)(\mathcal{T}_{K}g)(A):g\in\mathcal{H}^{0}\}$, and $\widehat{\mathcal{G}}=\{g(S)(\widehat{\mathcal{T}}_{K,0}g)(A):g\in\mathcal{G}^{0}\}$.
Let $\mathrm{star}(\mathcal{F})=\{\gamma g:\gamma\in[0,1],g\in\mathcal{F}\}$. Let $\widetilde{\delta}_{N}$, $\bar{\delta}_{N}$ and $\breve{\delta}_{N}$ be solutions to $\mathcal{R}_{N}(\mathrm{star}(\mathcal{F}),\delta)\leq\delta^{2}$ for $\mathcal{F}=\mathcal{G}^{0},\mathcal{G}^{1},\mathcal{T}_{K}\mathcal{G}$, respectively, let $\widetilde{\delta}_{n}$ be a solution to $\mathcal{R}_{n,a}(\mathrm{star}(\mathcal{G}^{0}),\delta)\leq\delta^{2}$ for every $a$, and let $\widehat{\delta}_{N}$ be a solution to $\mathcal{R}_{N}^{|0}(\mathrm{star}(\widehat{\mathcal{G}}),\delta)\leq\delta^{2}$. Let $\widehat{r}_{N}^{2}$ and $\bar{r}_{N}^{2}$ be bounds on $\mathcal{R}^{|0}(\widehat{\mathcal{G}},\infty)$ and $\mathcal{R}(\mathcal{G}^{1},\infty)$.
@@
$\|\mathcal{T}_{K}(\widehat{h}-h_{0})\|^{2}=O(\lambda+\delta_{N}\delta_{n})~{}~{}\text{with probability $1-O(\zeta)$, and}$
@@
3 An application of npJIVE: Confounding Robust Surrogate Indices from Many Weak Experiments
@@ 3.1 Background: Statistical-Surrogate Indices
Possibly the most common approach assumes that short-term observations form a statistical surrogate (Prentice, 1989). This assumes that the long-term outcome is conditionally independent of treatment given the short-term outcomes. This involves two key restrictions: that there are no unobserved confounding between short- and long-term outcomes (no $U_{1}$ in Fig. 2) and that all of the treatment’s effect on long-term outcomes is mediated by the short term (no $U_{2}$ in Fig. 2). As the latter restriction becomes more defensible as we include more short-term outcomes so as to mediate more of the treatment’s long-term effect, Athey et al. (2019) combine many short-term observations into a surrogate index, assuming they form a statistical surrogate and using historical data to regress long-term on short-term (or other ways of adjustment such as weighting). The proposal is simple, effective, and, as such, widely adopted.
However, even if short-term outcomes fully mediate the long-term effect, they may fail to satisfy statistical surrogacy. Consider the causal diagram in Fig. 2, where $S$ perfectly mediates $A$’s effect on $Y$ (i.e., exclusion restriction), but $S$ and $Y$ share an unobserved confounder $U$, while treatment $A$ is fully unconfounded. In this case, $S$ is a collider so that conditioning on it induces a path from $A$ to $Y$ via $U$, violating surrogacy and imperiling analysis using such methods as Athey et al. (2019).
@@ 3.2 Experiments as Instruments and Surrogates as Proxies
Athey et al. (2019) consider a setting where historical data prior to the present experiment contains only $S$ and $Y$ and some baseline covariates. In this context, all we can do is either worry about potential unobserved confounders, or hope the included covariates satisfy ignorability between $S$ and $Y$. However, at organizations that routinely run many digital experiments, we can take historical data from past experiments where we also observe the randomized treatments $A$. In the setting of Fig. 2, these treatments constitute an IV, which can help us identify the causal effect of $S$ on $Y$ and therefore infer the effect of a novel treatment on $Y$ by considering only its effect on $S$.
@@
In the following, we first present the corresponding causal model (summarized in LABEL:fig:complex). Then, we present an identification result showing how the long-term outcome of a novel treatment can be identified without observations on long-term outcomes under the novel treatment by leveraging historical experiments where we observe both short- and long-term outcomes. This identification result is written as a function of NPIV solutions, which may be estimated slowly in this nonparametric setting, imperiling rates on causal effect estimation if we were to simply plug-in an estimated solution. Therefore, we finally present a debiased formulation of the identification result that, under a strong identification condition, enjoys a mixed-bias property, which means we can multiply estimation rates and make them insignificant to first order. Together with estimation results for a nuisance in this new debiased formulation, we obtain asymptotic normal estimates for long-term causal effects of the novel treatment, enabling inference.
@@
Figure 2: A setting with unconfounded treatment ($A$) but confounded surrogate ($S$) and outcome ($Y$).
@@ 3.3 Causal Model
We state our causal model in general terms using potential outcomes. The model is summarized in Fig. 2, which represents a causal diagram that is consistent with these assumptions. We assume the existence of random variables $\widetilde{A},U,S(a),Y(s)$ and we refer to the variables $S(a)$ and $Y(s)$ as potential outcomes. We define $Y(a)=Y(S(a))$.
Here we let $\widetilde{A}\in\{1,\dots,K\}\cup\{\mathrm{new}\}$, where $\widetilde{A}\in\{1,\ldots,K\}$ denotes the historical experimental data and $\widetilde{A}=\mathrm{new}$ denotes the novel treatment (we will return to data generation after the causal model). We are interested in inference on the average outcome under the novel treatment:
$\theta_{0}=\mathbb{E}[Y(\mathrm{new})].$
@@
Assumption 1 (Structural relationship).
There exists a function $h^{\star}$ such that $\forall a\in\{0,\ldots,K\}$ $E[h^{\star}(S(a))]=E[Y(a)]$.
@@
$\widetilde{A}$ is independent of all potential outcomes and of $U$.
@@
It is straightforward to check that our causal assumptions hold if the potential outcomes are generated by the nonparametric structural causal model (SCM)
$Y=$ $f_{Y}(S,\epsilon_{Y})+g_{Y}(U,\epsilon_{Y})$
$S=$ $f_{S}(\widetilde{A},U,\epsilon_{S})$
$U=$ $f_{U}(\epsilon_{U})$
$\widetilde{A}=$ $f_{A}(\epsilon_{A})$
where $\epsilon_{Y}$, $\epsilon_{S}$, $\epsilon_{U}$, and $\epsilon_{A}$ are independent. The key for assumption 1 to hold is that $U$ and $S$ enter the outcome equation in an additively separable manner.
@@ 3.4 Data
Let $\widetilde{S}=S(\widetilde{A})$ and $\widetilde{Y}=Y(\widetilde{A})$ be the factually observed outcomes
@@
Let $S^{\mathrm{new}}\sim\widetilde{S}\mid\widetilde{A}=0$. We observe $n^{\prime}$ independent copies $S^{\mathrm{new}}_{1},\ldots,S^{\mathrm{new}}_{n^{\prime}}$ of $S^{\mathrm{new}}$.
We now turn to the question of identification. Note that had we observed $Y^{\mathrm{new}}\sim\widetilde{Y}\mid\widetilde{A}=0$ from the novel treatment, we would have trivially identified $\theta_{0}$ as $\mathbb{E}[Y^{\mathrm{new}}]$ under Assumption 3. So, in the absence of outcome observations for the novel treatment, our goal is to combine the historical and novel datasets to identify $\theta_{0}$.
@@ 3.5 Surrogacy and set identification
It is immediate under the structural relationship (assumption 1) that $h^{\star}(S)$ is a valid surrogate index.
That is, $h^{\star}$ connects causal effects on $Y$ to causal effects on $S$: if we could access $h^{\star}$, we would be able to know the average potential outcome of $Y$ under any treatment from the distribution of the average potential outcome of $S$ under that same intervention. We thus define the causal parameter of interest as $\theta^{\star}=\theta(h^{\star})$ where $\theta(h)=\mathbb{E}[h(S(\mathrm{new}))]$ for the novel intervention 00.
We consider a nuisance space $\mathcal{H}\subseteq L_{2}(\mathcal{S})$. Let $\mathcal{T}_{K}:\mathcal{H}\to([K]\to\mathbb{R})$ be the operator defined for every $h\in\mathcal{H}$ and every $a\in[K]$ by $(\mathcal{T}_{K}h)(a)=\mathbb{E}[h(S)\mid A=a]$.
@@
The following lemma characterizes $h^{\star}$ in terms of the data distribution in the historical cells.
@@
Under the causal assumptions 1-3, it holds that $h^{\star}\in\mathcal{H}_{0}$ where $\mathcal{H}_{0}$ is the set of functions $h:\mathcal{S}\to\mathbb{R}$ that satisfy the conditional moment restriction (CMR) $\mathbb{E}[h^{\star}(S)\mid A=a]=\mathbb{E}[Y\mid A=a]$ for every $a\in[K]$.
3.6 Identification and strong identification
While $h^{\star}$ is not point identified a priori, it is not our parameter of interest: $\theta(h^{\star})$ is. We provide below a formal definition of what it means for $\theta(h^{\star})$ to be identified.